A firm has been the product function Q= InL +2Ink. Derive the input demand law for both L and K in term of the product price P when the facor prices are PL =2, Ok =8. Hence derive the supply function for the firm product given the out as a function of the product price .

1
Expert's answer
2021-11-08T11:41:24-0500

Product function Q= InL +2Ink.


Deriving the input demand law for both L and K:


"\frac{\delta Q}{\delta L}=\frac{1}{L}+2lnK"


"\frac{\delta Q}{\delta K}=lnL+\frac{2}{K}"


Β Supply function for the firm product:


Basic form of a supply function is 𝑄=𝑄𝑆(𝑃)


Q=ln2+2ln8=4.85


"\therefore" 4.85=𝑄"_{s}" (𝑃)


Q"_{s}"="\frac{4.85}{P}"